# Ion Temperature Anisotropy Limits from Magnetic Curvature Scattering in Magnetotail Reconnection Jets

### v1

@@ Abstract
In collisionless plasmas, relaxation of the deviations of ion velocity distribution functions (VDFs) from local thermodynamic equilibrium (LTE) occurs through particle interactions with electromagnetic fields. In particular, in the Earth’s magnetotail, the deviations of the ion VDFs, typically consisting of multiple components, from the equilibrium must be limited to maintain stability of the current sheet. Curvature scattering is a leading candidate mechanism to limit such deviations, but its role remains insufficiently understood. We investigate the limits of ion temperature anisotropy in a magnetotail-like configuration by modeling a quasi-1D current sheet with a finite magnetic field curvature and three ion populations. We derive analytical thresholds for anisotropy based on current sheet stability and validate against spacecraft observations and numerical results. Our findings demonstrate that curvature scattering imposes limits on ion anisotropies, thereby maintaining the stability of the current sheet.
@@ I Introduction
Magnetic reconnection is a fundamental plasma process that explosively converts magnetic energy into particle kinetic and internal energy by reconfiguring the magnetic field topology. Magnetic reconnection powers energetic events in astrophysical systems such as black hole flares, accretion disks, and solar flares, plays a key role in magnetotail dynamics, and limits confinement in fusion plasmas [zweibel_magnetic_2009, yamada_magnetic_2010]. Extensive in situ spacecraft observations and numerical simulations demonstrated that multiple mechanisms conspire to energize charged particles (ions and electrons) during anti-parallel magnetic reconnection ($B_{g}=0$), as typically observed in the Earth’s magnetotail [norgren_electron_2025, oka_particle_2023]. In particular, ion energization primarily results from acceleration by the Hall and reconnection electric fields. In the reconnection inflow, ions are approximately in a local thermodynamic equilibrium characterized by a Maxwell-Boltzmann velocity distribution function (VDF). The Hall electric field associated with the quadrupolar magnetic field arising from electrons flowing toward and away from the X-line [uzdensky_physical_2006], ballistically accelerates ions from the reconnection inflow towards the center of the reconnecting current sheet, forming two cold beams counter-streaming along the normal magnetic field with parallel, i.e., magnetic field aligned, anisotropy [wygant_cluster_2005]. Alternatively, similar counter-streaming ion beams can form through pick-up acceleration, in which ions entering the reconnection outflow are swept up and accelerated by the bulk flow [drake_ion_2009, eastwood_ion_2015, liu_ionbeamdriven_2019]. Meanwhile, the reconnecting electric field accelerates demagnetized ions that undergo quasi-adiabatic transient Speiser orbits, characterized by a drifting, bean-shaped ion VDFs at the center of the current sheet [cowley_current_1983, burkhart_particle_1992]. Both mechanisms generate highly anisotropic ion VDFs that carry a substantial fraction of the current density.
@@
In situ observations in the Earth’s magnetotail demonstrated that the anisotropy of the ion VDFs is efficiently limited downstream in the reconnection outflow [wu_proton_2013, richard_fast_2023]. Typically, in collisionless plasmas, this relaxation is thought to be achieved through wave-particle interaction with electromagnetic waves excited by temperature anisotropy-driven instabilities [gary_theory_1993]. In contrast, within the reconnection outflow, numerical simulations and in situ observations indicate that the linear growth time of temperature–anisotropy–driven instabilities is small compared with the time it takes for the outflow to propagate across distances comparable to the system size [hietala_ion_2015, richard_fast_2023]. Consequently, the excited waves cannot reach large amplitudes. Therefore, pitch-angle scattering by wave-particle interaction cannot explain the fast relaxation of the ion VDFs into a lower energy state. Instead, in situ observations suggested that curvature scattering can provide such efficient relaxation [richard_fast_2023].
@@
In Earth’s magnetotail, in situ observations have identified three distinct populations in the ion VDFs corresponding to different regimes of motion [artemyev_ion_2019, kamaletdinov_ion_2025]. The first is a hot thermalized isotropic background. This population corresponds to the already-thermalized outflow. The second population is an adiabatic ($\kappa\gg 1$) population composed of two, cold, counter-streaming, field-aligned beams originating from ions ballistically accelerated by the Hall electric field in the reconnection region [wygant_cluster_2005], ions picked up by the outflow [drake_ion_2009, eastwood_ion_2015, liu_ionbeamdriven_2019], and ionospheric outflow [alm_magnetotail_2018, xu_ionospheric_2019, li_quantification_2021]. This parallel-anisotropic population contributes to the current density via curvature drift of its adiabatic orbits. The third population consists of supra-thermal ions accelerated by the cross-tail electric field $E_{y}$, driven by the reconnection electric field and by mesoscale convective structures such as bursty bulk flows [angelopoulos_statistical_1994]. These ions form a perpendicular, anisotropic, Speiser population, which contributes to the cross-tail current through their quasi-adiabatic motion ($\kappa\ll 1$). The interplay between the cold and Speiser anisotropic populations reflects a dynamic equilibrium within the current sheet, in which transient enhancements of anisotropy control the force balance, therefore controlling the stability of the current sheet [artemyev_ion_2019, an_configuration_2022, arnold_pic_2023]. This dynamical evolution of the interplay between ion anisotropy and the stability of the current sheet is an ongoing area of research. However, although ion anisotropy is recognized to be necessary to balance the magnetic tension [rich_balance_1972], the quantitative bounds of anisotropy that ensure current sheet stability are still not well constrained.
@@ II Theoretical Model
We consider a quasi-1D planar magnetic field configuration typical of the magnetotail current sheet [arnold_pic_2023, zhang_radial_2025], where the reconnecting component of the magnetic field satisfies $\partial_{z}B_{x}\simeq\mu_{0}j_{y}$, the normal component is approximately constant ($B_{z}\simeq\mathrm{const}$), and the curvature radius of the magnetic field is $R_{c}\simeq B_{z}/\mu_{0}j_{y}$. These current sheets are often polarized by strong electric fields $E_{z}$ that redistribute ion and electron currents through $\bm{E}\times\bm{B}$ drifts [hesse_ionscale_1998, lu_hall_2016]. However, within the reference frame moving with the $\bm{E}\times\bm{B}$ drift, the electron current is primarily dominated by diamagnetic drifts [birn_thin_2004] and curvature drifts [artemyev_contribution_2019].
@@ II.1 Current Density and Adiabaticity Parameter
To estimate the ion contribution to the current density, we model the ion VDFs as consisting of three primary populations as identified in the Earth’s magnetotail [artemyev_ion_2019, kamaletdinov_ion_2025]: (i) a cold ion population ($\kappa>1$) composed of two field-aligned counter-streaming beams with number density $n_{c}$ and temperature $T_{c}=(T_{c\parallel}+2T_{c\bot})/3$, with a parallel anisotropy $A_{c}=T_{c\parallel}/T_{c\bot}>1$; (ii) a hot, nearly isotropic, population with density $n_{h}$ and temperature $T_{h}\approx T_{h\parallel}\approx T_{h\bot}$; (iii) an energetic Speiser-type population ($\kappa\ll 1$) with density $n_{s}$, temperature $T_{s}=(T_{s\parallel}+2T_{s\bot})/3$, and perpendicular anisotropy $A_{s}=T_{s\parallel}/T_{s\bot}<1$. We note that, as we will show later in Sections III and IV, the relative contribution of each population can vary across the current sheet [hoshino_ion_1998, hietala_ion_2015, kamaletdinov_ion_2025].
Given the typical relations $n_{s}\ll n_{c}\sim n_{h}$ and $T_{c}\ll T_{h}<T_{s}$ [artemyev_ion_2019], we can approximate the total ion number density as $n_{i}=n_{c}+n_{h}+n_{s}\approx n_{c}+n_{h}$, the parallel and perpendicular ion pressures as $P_{i\parallel,\bot}=n_{i}k_{B}T_{i\parallel,\bot}\approx P_{h}+P_{s\parallel,\bot}$, and the ion anisotropy as
$A_{i}-1=\dfrac{T_{i\parallel}}{T_{i\bot}}-1\approx\dfrac{n_{c}}{n_{i}}\left(A_{c}-1\right)+\dfrac{n_{s}}{n_{i}}\left(A_{s}-1\right).$ (1)
The different ion populations contribute to the total current density through different drifts. Due to its parallel anisotropy and low thermal energy, the cold ion population primarily contributes via anisotropic curvature drifts [krall_principles_1973]:
$j_{cy}\approx\dfrac{n_{c}k_{B}T_{c\bot}(A_{c}-1)}{B^{2}}\dfrac{\partial B_{x}}{\partial z}=\dfrac{n_{c}}{n_{i}}(A_{c}-1)\eta_{c}^{2}\dfrac{\beta_{i\parallel}}{2}j_{y},$ (2)
@@
Due to its large temperature and isotropy, the hot ion population primarily contributes via diamagnetic drifts. This gives $j_{hy}\simeq\left[B_{x}\partial_{z}P_{h}-B_{z}\partial_{x}P_{h}\right]/B^{2}$. Thus, in current sheets with finite $B_{z}$, this term reduces to $j_{hy}\simeq-\partial_{x}P_{h}/B_{z}$ around the equatorial plane ($B_{x}\sim 0$) and becomes significant only in quasi-2D configurations where $\partial_{x}P_{h}/B_{z}$ is comparable to $\partial_{z}P_{h}/B_{x}$ [birn_thin_2004]. Although such quasi-2D current sheets can be found in the magnetotail [apatenkov_multispacecraft_2025], thin, intense current sheets generally show very weak $\partial_{x}P_{h}$ gradient, such that $\partial_{x}P_{h}/B_{z}\ll\partial_{z}P_{h}/B_{x}$ [artemyev_configuration_2021, zhang_radial_2025]. Therefore, in such thin, intense current sheets, the hot ion contribution to the cross-tail current is negligible $j_{hy}\approx 0$ around the equatorial plane.
Speiser ions contribute to the current density through their quasi-adiabatic trajectories. Using an approximation of the Speiser ion VDF [burkhart_ion_1992, burkhart_particle_1992, artemyev_proton_2010], the current density from this population can be estimated as $j_{sy}\simeq en_{s}v_{sD}$, where $v_{sD}$ is the drift speed of the Speiser population. Defining a characteristic current $j_{0}=env_{iT\bot}$ where $v_{iT\bot}=\sqrt{2k_{B}T_{i\bot}/m_{i}}$ is the perpendicular thermal speed, we express the Speiser ion contribution to the current density as $j_{sy}\simeq(n_{s}/n)\eta_{s}j_{0}$, with $\eta_{s}=v_{sD}/v_{iT\bot}$.
Summing these components plus the electron current density $j_{ey}=\eta_{e}j_{0}$, where $\eta_{e}$ is the fraction of electron to total current, the total current density, $j_{y}=j_{ey}+j_{cy}+j_{hy}+j_{sy}$, can be written as
$j_{y}=j_{0}\dfrac{{\eta_{e}+\dfrac{n_{s}}{n_{i}}\eta_{s}}}{{1-\dfrac{n_{c}}{n_{i}}(A_{c}-1)\eta_{c}^{2}\dfrac{\beta_{i\parallel}}{2}}}.$ (3)
@@
$\kappa^{2}=\dfrac{1}{\beta_{i\parallel}}\dfrac{T_{i\parallel}}{T_{i\bot}}\dfrac{1-\dfrac{n_{c}}{n_{i}}(A_{c}-1)\eta_{c}^{2}\dfrac{\beta_{i\parallel}}{2}}{{\eta_{e}+\dfrac{{n_{s}}}{n_{i}}\eta_{s}}}.$ (4)
In the limit of isotropic ions ($n_{s}/n_{i}=n_{c}/n_{i}\ll 1$, $T_{i\parallel}/T_{i\bot}=1$), we obtain $j_{y}=\eta_{e}j_{0}$, and $\kappa^{2}=1/\eta_{e}\beta_{i}=d_{i}^{2}/\eta_{e}\rho_{i}^{2}$, i.e., $R_{c}=d_{i}^{2}/\eta_{e}\rho_{i}$, where $d_{i}=\sqrt{m_{i}/\mu_{0}n_{0}e^{2}}$ is the ion inertial length. Below, we investigate the limits of parallel and perpendicular ion anisotropy, so that $j_{y}$ remains limited, ensuring stability of the current sheet.
@@ II.2.1 Cold ions dominated regime
In the first case, we consider that the ion population consists mainly of hot isotropic and cold ions, whereas Speiser ions are negligible ($n_{s}\approx 0$, $j_{sy}\approx 0$). Since cold ions are field-aligned counter-streaming beams, this regime corresponds to a parallel anisotropy, and Eq. 1 becomes
$\dfrac{n_{c}}{n_{i}}\left(A_{c}-1\right)=\left(\dfrac{T_{i\parallel}}{T_{i\bot}}-1\right)\geq 0.$ (5)
Thus, the current density Eq. 3 becomes
$j_{y}=j_{0}\dfrac{\eta_{e}}{{1-\left(\dfrac{T_{i\parallel}}{T_{i\bot}}-1\right)\eta_{c}^{2}\dfrac{\beta_{i\parallel}}{2}}},$ (6)
@@
$\kappa^{2}=\dfrac{1}{\beta_{i\parallel}}\dfrac{T_{i\parallel}}{T_{i\bot}}\dfrac{1}{\eta_{e}}\left[1-\left(\dfrac{T_{i\parallel}}{T_{i\bot}}-1\right)\eta_{c}^{2}\dfrac{\beta_{i\parallel}}{2}\right].$ (7)
@@
$\mathcal{R}_{c}(\eta_{c},\beta_{i\parallel})=\left(1+\dfrac{2}{{\eta_{c}^{2}\beta_{i\parallel}}}\right)^{-1},$ (8)
where $\mathcal{R}_{c}(\eta_{c},\beta_{i\parallel})$ denotes the anisotropy at the stability threshold. In the limit $\eta_{c}^{2}=T_{c\bot}/T_{i\parallel}=1$, we recover the classic fluid firehose stability condition where the ion inertia balances the magnetic tension force [cowley_plasma_1980].
Figure 1: (a) Current density and (b) adiabaticity parameter in the model 1D current sheet with three ion populations with $\eta_{e}=$, $\eta_{s}=$, and $\eta_{c}=$. The parallel anisotropy limit [Eqs. 6 and 7] is shown in the lower half of each panel $T_{i\bot}/T_{i\parallel}<1$, and the perpendicular anisotropy limit [Eqs. 10 and 11] is shown in the upper half $T_{i\bot}/T_{i\parallel}>1$. The black thick line indicate $\kappa=1$ and the black thin lines indicate $\kappa=10^{-2},\,10^{-1},\,10^{1},\,10^{2}$. The dashed cyan lines indicate the anisotropy thresholds [Eqs. 8 and 12].

### v2

@@ Abstract
In collisionless plasmas, relaxation of the deviations of ion velocity distribution functions (VDFs) from local thermodynamic equilibrium occurs through particle interactions with electromagnetic fields. In particular, in the Earth’s magnetotail, the deviations of the ion VDFs, typically consisting of multiple components, from the equilibrium must be limited to maintain stability of the current sheet. Curvature scattering is a leading candidate mechanism to limit such deviations, but its role remains insufficiently understood. We investigate the limits of ion temperature anisotropy in a magnetotail-like configuration by modeling a quasi-1D current sheet with a finite magnetic field curvature and three ion populations. We derive analytical thresholds for anisotropy based on current sheet stability and validate against spacecraft observations and numerical simulations. Our findings demonstrate that curvature scattering imposes limits on ion anisotropies, thereby maintaining the stability of the current sheet.
@@ I Introduction
Magnetic reconnection is a fundamental plasma process that explosively converts magnetic energy into particle kinetic and internal energy by reconfiguring the magnetic field topology. Magnetic reconnection powers energetic events in astrophysical systems such as black hole flares, accretion disks, and solar flares, plays a key role in magnetotail dynamics, and limits confinement in fusion plasmas [82, 77]. Extensive in situ spacecraft observations and numerical simulations demonstrated that multiple mechanisms conspire to energize charged particles (ions and electrons) during anti-parallel magnetic reconnection ($B_{g}=0$), as typically observed in the Earth’s magnetotail [54, 56]. In particular, ion energization primarily results from acceleration by the Hall and reconnection electric fields. In the reconnection inflow, ions are approximately in a local thermodynamic equilibrium characterized by a Maxwell-Boltzmann velocity distribution function (VDF). The Hall electric field associated with the quadrupolar magnetic field arising from electrons flowing toward and away from the X-line [72], ballistically accelerates ions from the reconnection inflow towards the center of the reconnecting current sheet, forming two cold beams counter-streaming along the normal magnetic field with parallel, i.e., magnetic field aligned, anisotropy [75]. Alternatively, similar counter-streaming ion beams can form through pick-up acceleration, in which ions entering the reconnection outflow are swept up and accelerated by the bulk flow [32, 33, 48]. Meanwhile, the reconnecting electric field accelerates demagnetized ions that undergo quasi-adiabatic transient Speiser orbits, characterized by a drifting, bean-shaped ion VDFs at the center of the current sheet [26, 19]. Both mechanisms generate highly anisotropic ion VDFs that carry a substantial fraction of the current density.
@@
In situ observations in the Earth’s magnetotail demonstrated that the anisotropy of the ion VDFs is efficiently limited downstream in the reconnection outflow [74, 65]. Typically, in collisionless plasmas, this relaxation is thought to be achieved through wave-particle interaction with electromagnetic waves excited by temperature anisotropy-driven instabilities [35]. In contrast, within the reconnection outflow, numerical simulations and in situ observations indicate that the linear growth time of temperature–anisotropy–driven instabilities is small compared with the time it takes for the outflow to propagate across distances comparable to the system size [40, 65]. Consequently, the excited waves cannot reach large amplitudes. Therefore, pitch-angle scattering by wave-particle interaction cannot explain the fast relaxation of the ion VDFs into a lower energy state. Instead, in situ observations suggested that curvature scattering can provide such efficient relaxation [65].
@@
In situ observations in the Earth’s magnetotail [8, 43] and numerical simulations of magnetotail reconnection [41, 40] have identified three distinct populations in the ion VDFs corresponding to different regimes of motion in the reconnection outflow. The first is a hot thermalized isotropic background. This population corresponds to the already-thermalized outflow. The second population is an adiabatic ($\kappa\gg 1$) population composed of two, cold, counter-streaming, field-aligned beams originating from ions ballistically accelerated by the Hall electric field in the reconnection region [75], ions picked up by the outflow [32, 33, 48], and ionospheric outflow [2, 76, 47]. This parallel-anisotropic population contributes to the current density via anisotropic curvature drift. The third population consists of supra-thermal ions accelerated by the cross-tail electric field $E_{y}$, driven by the reconnection electric field and by mesoscale convective structures such as bursty bulk flows [64]. These ions form a perpendicular, anisotropic, Speiser population, which contributes to the cross-tail current through their quasi-adiabatic motion ($\kappa\ll 1$). The interplay between the cold and Speiser anisotropic populations reflects a dynamic equilibrium within the current sheet, in which transient enhancements of anisotropy control the force balance, therefore controlling the stability of the current sheet [8, 3, 6]. This dynamical evolution of the interplay between ion anisotropy and the stability of the current sheet is an ongoing area of research. However, although ion anisotropy is recognized to be necessary to balance the magnetic tension [61], the quantitative bounds of anisotropy that ensure current sheet stability are still not well constrained.
@@ II Theoretical Model
We consider a quasi-1D planar magnetic field configuration typical of the magnetotail current sheet [6, 81], where the reconnecting component of the magnetic field satisfies $\partial_{z}B_{x}\simeq\mu_{0}j_{y}$, the normal component is approximately constant ($B_{z}\simeq\mathrm{const}$), and the curvature radius of the magnetic field is $R_{c}\simeq B_{z}/\mu_{0}j_{y}$. These current sheets are often polarized by strong electric fields $E_{z}$ that redistribute ion and electron currents through $\bm{E}\times\bm{B}$ drifts [39, 51]. However, within the reference frame moving with the $\bm{E}\times\bm{B}$ drift, the electron current is primarily dominated by diamagnetic drifts [16] and curvature drifts [7].
@@ II.1 Current Density and Adiabaticity Parameter
To estimate the ion contribution to the current density, we model the ion VDFs as consisting of three primary populations as identified in the Earth’s magnetotail [8, 43]: (i) a cold ion population ($\kappa>1$) composed of two field-aligned counter-streaming beams with number density $n_{c}$ and temperature $T_{c}=(T_{c\parallel}+2T_{c\bot})/3$, with a parallel anisotropy $A_{c}=T_{c\parallel}/T_{c\bot}>1$; (ii) a hot, nearly isotropic, population with density $n_{h}$ and temperature $T_{h}\approx T_{h\parallel}\approx T_{h\bot}$; (iii) an energetic Speiser-type population ($\kappa\ll 1$) with density $n_{s}$, temperature $T_{s}=(T_{s\parallel}+2T_{s\bot})/3$, and perpendicular anisotropy $A_{s}=T_{s\parallel}/T_{s\bot}<1$. Consequently, cold ions mainly contribute to the parallel temperature anisotropy of the superposed three species population, whereas Speiser ions are responsible for the perpendicular anisotropy and agyrotropy through their quasi-adiabatic, unmagnetized orbits. We note that, as we will show later in Sections III and IV, the relative contribution of each population can vary across the current sheet [41, 40, 43].
The different ion populations contribute to the total current density through different drifts. Due to its parallel anisotropy and low thermal energy, the cold ion population primarily contributes via anisotropic curvature drifts [44]:
[…]
@@
Due to its large temperature and isotropy, the hot ion population primarily contributes via diamagnetic drifts. This gives $j_{hy}\simeq\left[B_{x}\partial_{z}P_{h}-B_{z}\partial_{x}P_{h}\right]/B^{2}$. Thus, in current sheets with finite $B_{z}$, this term reduces to $j_{hy}\simeq-\partial_{x}P_{h}/B_{z}$ around the equatorial plane ($B_{x}\sim 0$) and becomes significant only in quasi-2D configurations where $\partial_{x}P_{h}/B_{z}$ is comparable to $\partial_{z}P_{h}/B_{x}$ [16]. Although such quasi-2D current sheets can be found in the magnetotail [5], thin, intense current sheets generally show very weak $\partial_{x}P_{h}$ gradient, such that $\partial_{x}P_{h}/B_{z}\ll\partial_{z}P_{h}/B_{x}$ [12, 81]. Therefore, in such thin, intense current sheets, the hot ion contribution to the cross-tail current is negligible $j_{hy}\approx 0$ around the equatorial plane.
Speiser ions contribute to the current density through their quasi-adiabatic trajectories. Using an approximation of the Speiser ion VDF [20, 19, 10], the current density from this population can be estimated as $j_{sy}\simeq en_{s}v_{sD}$, where $v_{sD}$ is the drift speed of the Speiser population. Defining a characteristic current $j_{0}=en_{i}v_{iT\bot}$ where $v_{iT\bot}=\sqrt{2k_{B}T_{i\bot}/m_{i}}$ is the perpendicular thermal speed, we express the Speiser ion contribution to the current density as $j_{sy}\simeq(n_{s}/n_{i})\eta_{s}j_{0}$, with $\eta_{s}=v_{sD}/v_{iT\bot}$.
Summing these components plus the electron current density $j_{ey}=\eta_{e}j_{0}$, where $\eta_{e}$ is the fraction of electron to characteristic current, the total current density, $j_{y}=j_{ey}+j_{cy}+j_{hy}+j_{sy}$, can be written as
[…]
@@
[…]
In the limit of isotropic ions ($n_{s}/n_{i}=n_{c}/n_{i}\ll 1$, $T_{i\parallel}/T_{i\bot}=1$), we obtain $j_{y}=\eta_{e}j_{0}$, and $\kappa^{2}=1/\eta_{e}\beta_{i}=d_{i}^{2}/\eta_{e}\rho_{i}^{2}$, i.e., $R_{c}=d_{i}^{2}/\eta_{e}\rho_{i}$, where $d_{i}=\sqrt{m_{i}/\mu_{0}n_{i}e^{2}}$ is the ion inertial length. Below, we investigate the limits of parallel and perpendicular ion anisotropy, so that $j_{y}$ remains limited, ensuring stability of the current sheet.
@@ II.2.1 Cold ions dominated regime
In the first case, we consider that the ion population consists mainly of hot isotropic and cold ions, whereas Speiser ions are negligible ($n_{s}\approx 0$, $j_{sy}\approx 0$). Since cold ions are field-aligned counter-streaming beams, this regime corresponds to a parallel anisotropy with
[…]
where we have assumed $T_{i\bot}\approx T_{c\bot}$. Thus, the current density Eq. 2 becomes
[…]
@@
[…]
@@
[…]
where $\mathcal{R}_{c}(\eta_{c},\beta_{i\parallel})$ denotes the anisotropy at the stability threshold. In the limit $\eta_{c}^{2}=T_{c\bot}/T_{i\parallel}=1$, we recover the classic fluid firehose stability condition where the ion inertia balances the magnetic tension force [25].
@@
Figure 1: (a) Adiabaticity parameter and (b) current density in the model 1D current sheet with three ion populations with $\eta_{e}=0.3$, $\eta_{s}=0.2$, and $\eta_{c}=3$. The parallel anisotropy limit [Eqs. 5 and 6] is shown in the lower half of each panel $T_{i\bot}/T_{i\parallel}<1$, and the perpendicular anisotropy limit [Eqs. 9 and 10] is shown in the upper half $T_{i\bot}/T_{i\parallel}>1$. The black thick line indicate $\kappa=1$ and the black thin lines indicate $\kappa=10^{-2},\,10^{-1},\,10^{1},\,10^{2}$. The dashed cyan and pink lines indicate the anisotropy thresholds [Eqs. 7 and 11].